# Switching the Loss Reduces the Cost in Offline Reinforcement Learning

### 08 Mar 2024 v1

@@ 1 Introduction
The intuition behind training FQI with log-loss comes from the observation that not all state-action-to-next-state transitions are equally noisy. Whereas training with squared loss amounts to assuming the noise is homoskedastic across transitions, training with log-loss learns to emphasize less noisy transitions and deemphasize more noisy transitions. Since in RL typically some transitions are noisier than others, learning to emphasize less noisy transitions leads to more sample efficient learning.
@@ 3 FQI-log: Fitted Q-Iteration with log-loss
The differences between log-loss and squared loss are made apparent in regression. As is well known, the ideal loss in regression compensates when the outcomes of the observed data points have unequal variances, a condition often called “heteroscedastic regression”. In heteroscedastic regression, it is desirable to deemphasize noisy outcomes, i.e. outcomes whose noise has high variance, and emphasize less noisy outcomes. As opposed to squared loss, which treats each data point as having equal-variance outcomes (“homoscedastic regression”), log-loss assigns a higher cost size=,color=green!20!white,]AA: Very ambiguous? What is cost here. Cs: cost is what you pay. I think cost work better here. to any predictor, $f$, that is closer to zero, which is obvious from its definition in Eq. 2. size=,color=orange!20!white,]SR: Should this say “to any predictor that is closer to 0 or 1”? This makes sense: if the response is nonnegative valued and bounded, predicting that its mean is small implies that the variance of the response is also small, hence, higher cost should be incurred when a small value is predicted.
@@ 4.1 Assumptions
We introduce the necessary assumptions under which our main theorem holds. size=,color=red!20!white,inline]Cs: I added the word nonstationary here. Discuss etc. Nonstationary policies need to be introduced. size=,color=green!20!white,inline]AA: I added nonstationary policies to the preliminaries section. right after we introduct stationary policies.
@@
There exists $B>0$ such that, for all $f\in\mathcal{F}$ and $i\in[n]$, $(\mathcal{T}f)(S_{i},A_{i})$ and $C_{i}+\gamma f^{\wedge}(S_{i}^{\prime})$ are in $[0,B]$ almost surely. size=,color=orange!20!white,]SR: This was a little tricky to read I think, $\mathcal{T}f$ is $\mu$-almost surely but it’s not super obvious that’s what’s meant. I changed to $(\mathcal{T}f)(S_{i},A_{i})$. Also I switched the quantification around, there exists $B$ such that for all … instead of for all … there exists $B$. I’m pretty sure the first version wasn’t what we need. size=,color=orange!20!white,]SR: Also we don’t really need this to hold for all $i$, if it holds for one it will hold for all since the data are iid and it’s a statement about probabilities.
@@ 5 Numerical Experiments
The goal of our experiments is to provide insights into the benefits of using FQI-log for learning a near-optimal policy in batch reinforcement learning. We run our first set of experiments in reinforcement learning with linear function approximation, as this setting allows us to best compare FQI-log to FQI-sq without other confounding factors. We then run experiments in deep reinforcement learning to verify whether our findings generalize to more complex function classes. The objective is either to minimize squared loss, which does not get small-cost bounds, or to minimize size=,color=green!20!white,]AA: binar log-loss, as proposed in this work.
@@ 5.2 Mountain Car
We first evaluate FQI-log and FQI-sq on an episodic sparse cost variant of Mountain Car (Moore, 1990), where the value functions are indexed by the timestep, i.e. $v_{h}$. size=,color=orange!20!white,]SR: This is just finite horizon right? The initialization is fixed so it’s not really episodic imo This environment consists of a $2$-dimensional continuous state space and $3$ discrete actions. The agent, i.e. the policy $\pi_{f_{k}}$, starts from a fixed state, $s_{1}=[-0.5,0]$, and interacts with the environment for 800 timesteps before an episode terminates. If an agent reaches the top of the hill, it remains there until the end of the episode. The cost is 0 at all timesteps except the last, when a cost of 1 is received if the agent has not reached the hilltop, and otherwise the cost is 0. A discount factor of $\gamma=1$ is used in our experiments.
@@ 7 Conclusions
size=,color=red!20!white,inline]Cs: What did we learn? Log-loss is good. But we need problems with small costs. Is this good? When? Goal-oriented stuff. What remains to be done? Undiscounted? Online?
W showed that in batch RL the loss function genuinely matters. We proved that FQI-log is more sample efficient on MDPs where the optimal cost $\bar{v}^{\star}$ is small. FQI-log is the first efficient batch RL algorithm to get a small-cost bound.
@@ A.2 Concentration for the log-loss estimator
Given a dataset $D=\{(S_{i},A_{i},C_{i},S_{i}^{\prime})\}_{i=1}^{n}$ with $n\in\mathbb{N}$ and a finite function class $\mathcal{F}\subseteq(\mathcal{S}\times\mathcal{A}\rightarrow[0,1])$ that satisfy Assumptions 4.2, 4.3, 4.5 and 4.6, $B=1$, it holds with probability $1-\delta$, the output policy of FQI-log after $k$ iterations, $\pi_{k}=\pi_{f_{k}}$, satisfies size=,color=red!20!white,]Cs: state this with some universal constant. do not startle the reviewer with big constants. explain somewhere that the constants are unoptimized..? could they be smaller..? etc.
@@
Since $\eta_{h}^{k}\times\pi_{k}$ and $\eta_{h}^{k}\times\pi^{\star}$ are admissible, as can be easily seen with an argument similar to that used in the proof of Lemma B.16, size=,color=red!20!white,]Cs: ok, this is out of sequence. figure out the best way to resolve this. maybe after the definition of admissibility add a little lemma on this and just use that here and also in the proof of Lemma B.16. it follows that with probability $1-\delta$,
@@
For policies $\pi,\bar{\pi}:\mathcal{S}\rightarrow\mathcal{M}_{1}(\mathcal{A})$, we have size=,color=red!20!white,]Cs: so if we use the convention that distributions of row vectors, we can also write $\left\langle\eta_{h}^{\pi},\,q^{\pi^{\prime}}(\cdot,\pi)-v^{\tilde{\pi}}\right\rangle$ as $\eta_{h}^{\pi}(q^{\tilde{\pi}}(\cdot,\pi)-v^{\tilde{\pi}})$. But I was wondering whether inner products have been introduced. size=,color=green!20!white,]AA: It is defined in the preliminaries.
@@
size=,color=red!20!white,inline]Cs: (defn of $v^{\star}$ or properties of $v^{\star}/q^{\star}$. By the way, these should be mentioned upfront! Maybe in the notation section and we can refer back to there. Now, size=,color=red!20!white,]Cs: learn to use MoveEqLeft, see next line in the source. Also learn to use qedhere.
@@
Let $f:\mathcal{S}\times\mathcal{A}\rightarrow[0,\infty)$ and let $\pi=\pi_{f}$ be a policy that is greedy with respect to $f$. size=,color=red!20!white,]Cs: inspired by lemma 1 byDylan Define $D_{f}=\sup_{h\geq 1}\max(\left\|\Delta_{f}\right\|_{2,\eta_{h}^{\pi}\times\pi},\left\|\Delta_{f}\right\|_{2,\eta_{h}^{\pi}\times\pi^{\star}})$. Then, it holds that
@@
As explained earlier, the analysis in this section uses contraction arguments that have a long history in the analysis of dynamic programming algorithms in the context of MDPs. size=,color=red!20!white,]Cs: Bertsekas’ book a good start, Howard, Whittle, etc. The novelty is that we need to bound the triangular deviation to $q^{*}$. As this has been shown to be upper bounded by the Hellinger distance (Corollary A.2), we switch to Hellinger distances and establishes contraction properties of $\mathcal{T}$ with respect to such distances. This required new proofs. The change of measure arguments used in the “error propagation analysis” are standard. size=,color=red!20!white,]Cs: need to add citations; Remi’s first paper, our FQI paper, Nan’s paper.
The following lemma, at a high level, establishes that the map $f\mapsto f^{\wedge}$ (“$\min$-operator”) is a non-expansion over the set of nonnegative functions with domain $\mathcal{S}\times\mathcal{A}$ and $\mathcal{S}$, respectively, when these function spaces are equipped with appropriate norms: size=,color=red!20!white,]Cs: I leave these references here. Are they relevant? Perhaps [Chen and Jiang, 2019] is? Why cite the others? [tsitsiklis1996feature, Littman, 1996]. Do we want to remind the reader that the $\min$-operator (or, equivalently, the $\max$) is a non-expansion for max-norm? and that this is at the heart of the analysis? maybe this is not the best place to do it and also the max-norm analysis is not super relevant here..
@@
Now, since $f$ is continuous over its domain, by the mean-value theorem, $f$ is decreasing over $[0,\infty)$. ∎size=,color=green!20!white,]AA: We can also use the algebriac proof given below. size=,color=red!20!white,]Cs: Note that if $b=0$, $a=1$, for example, the right-side derivative at $x=0$ blows up; $f$ is not differentiable on the right etc.!
@@
Let $\mu,\nu$ be any distributions over $\mathcal{S}\times\mathcal{A}$ and assume that $\nu$ is admissible. size=,color=red!20!white,]Cs: somewhere we need to say that in $\nu P\times\pi$, the multiplication between $\nu$ and $P$ has a higher precedence than $\times$. Or put parentheses everywhere. Then, for any $f,f^{\prime}:S\times A\rightarrow[0,\infty)$ we have
@@
Define $(\nu_{i})_{0\leq i\leq k}$ via $\nu_{k}=\nu$ and for $0\leq i<k$, let $\nu_{i}=(\nu_{i+1}P)\times\pi_{f_{i},q^{*}}$. size=,color=red!20!white,]Cs: if we clean up $\nu P\times\pi$, the parentheses here and later could be removed. Note that by assumption, $\nu_{k}$ is admissible. It then follows that $\nu_{i}$ for $0\leq i<k$ is also admissible. Indeed, if for some $0\leq i<k$, $\pi=(\pi_{0},\pi_{1},\dots)$ is the nonstationary policy that realizes $\nu_{i+1}$ in step $s\geq 0$, $\pi^{\prime}=(\pi_{0},\pi_{1},\dots,\pi_{s},\pi_{f_{i},q^{*}},\pi_{s+1},\dots)$ is a policy that realizes $\nu_{i}$ in step $s+1$.
@@
Now, we bound $S$ defined above: size=,color=red!20!white,]Cs: to save a parenthesis, I use that operator application takes precedence over function evaluation

### 11 Mar 2024 v2

@@ 1 Introduction
The intuition behind training FQI with log-loss comes from the observation that not all state-action-to-next-state transitions are equally noisy. Whereas training with squared loss amounts to assuming the noise is homoscedastic across transitions, training with log-loss learns to emphasize less noisy transitions and de-emphasize more noisy transitions. Since in RL typically some transitions are noisier than others, learning to emphasize less noisy transitions leads to more sample efficient learning.
@@ 3 FQI-log: Fitted Q-Iteration with log-loss
The differences between log-loss and squared loss are made apparent in regression. As is well known, the ideal loss in regression compensates when the outcomes of the observed data points have unequal variances, a condition often called “heteroscedastic regression”. In heteroscedastic regression, it is desirable to de-emphasize noisy outcomes, i.e. outcomes whose noise has high variance, and emphasize less noisy outcomes. As opposed to squared loss, which treats each data point as having equal-variance outcomes (“homoscedastic regression”), log-loss assigns a higher cost size=,color=green!20!white,]AA: Very ambiguous? What is cost here. Cs: cost is what you pay. I think cost work better here. to any predictor, $f$, that is closer to zero, which is obvious from its definition in Eq. 2. size=,color=orange!20!white,]SR: Should this say “to any predictor that is closer to 0 or 1”? This makes sense: if the response is nonnegative valued and bounded, predicting that its mean is small implies that the variance of the response is also small, hence, higher cost should be incurred when a small value is predicted.
@@ 4.1 Assumptions
We introduce the necessary assumptions under which our main theorem holds.
@@
There exists $B>0$ such that, for all $f\in\mathcal{F}$ and $i\in[n]$, $(\mathcal{T}f)(S_{i},A_{i})$ and $C_{i}+\gamma f^{\wedge}(S_{i}^{\prime})$ are in $[0,B]$ almost surely. size=,color=orange!20!white,]SR: This was a little tricky to read I think, $\mathcal{T}f$ is $\mu$-almost surely but it’s not super obvious that’s what’s meant. I changed to $(\mathcal{T}f)(S_{i},A_{i})$. Also I switched the quantification around, there exists $B$ such that for all … instead of for all … there exists $B$. I’m pretty sure the first version wasn’t what we need.
@@ 5 Numerical Experiments
The goal of our experiments is to provide insights into the benefits of using FQI-log for learning a near-optimal policy in batch reinforcement learning. We run our first set of experiments in reinforcement learning with linear function approximation, as this setting allows us to best compare FQI-log to FQI-sq without other confounding factors. We then run experiments in deep reinforcement learning to verify whether our findings generalize to more complex function classes. The objective is either to minimize squared loss, which does not get small-cost bounds, or to minimize log-loss, as proposed in this work.
@@ 5.2 Mountain Car
We first evaluate FQI-log and FQI-sq on an episodic sparse cost variant of Mountain Car (Moore, 1990), where the value functions are indexed by the timestep, i.e. $v_{h}$. This environment consists of a $2$-dimensional continuous state space and $3$ discrete actions. The agent, i.e. the policy $\pi_{f_{k}}$, starts from a fixed state, $s_{1}=[-0.5,0]$, and interacts with the environment for 800 timesteps before an episode terminates. If an agent reaches the top of the hill, it remains there until the end of the episode. The cost is 0 at all timesteps except the last, when a cost of 1 is received if the agent has not reached the hilltop, and otherwise the cost is 0. A discount factor of $\gamma=1$ is used in our experiments.
@@ 7 Conclusions
We showed that in batch RL the loss function genuinely matters. We proved that FQI-log is more sample efficient on MDPs where the optimal cost $\bar{v}^{\star}$ is small. FQI-log is the first efficient batch RL algorithm to get a small-cost bound.
@@ A.2 Concentration for the log-loss estimator
Given a dataset $D=\{(S_{i},A_{i},C_{i},S_{i}^{\prime})\}_{i=1}^{n}$ with $n\in\mathbb{N}$ and a finite function class $\mathcal{F}\subseteq(\mathcal{S}\times\mathcal{A}\rightarrow[0,1])$ that satisfy Assumptions 4.2, 4.3, 4.5 and 4.6, $B=1$, it holds with probability $1-\delta$, the output policy of FQI-log after $k$ iterations, $\pi_{k}=\pi_{f_{k}}$, satisfies
@@
Since $\eta_{h}^{k}\times\pi_{k}$ and $\eta_{h}^{k}\times\pi^{\star}$ are admissible, as can be easily seen with an argument similar to that used in the proof of Lemma B.16, it follows that with probability $1-\delta$,
@@
For policies $\pi,\bar{\pi}:\mathcal{S}\rightarrow\mathcal{M}_{1}(\mathcal{A})$, we have
@@
size=,color=red!20!white,inline]Cs: (defn of $v^{\star}$ or properties of $v^{\star}/q^{\star}$. By the way, these should be mentioned upfront! Maybe in the notation section and we can refer back to there. Now,
@@
As explained earlier, the analysis in this section uses contraction arguments that have a long history in the analysis of dynamic programming algorithms in the context of MDPs. The novelty is that we need to bound the triangular deviation to $q^{*}$. As this has been shown to be upper bounded by the Hellinger distance (Corollary A.2), we switch to Hellinger distances and establishes contraction properties of $\mathcal{T}$ with respect to such distances. This required new proofs. The change of measure arguments used in the “error propagation analysis” are standard.
The following lemma, at a high level, establishes that the map $f\mapsto f^{\wedge}$ (“$\min$-operator”) is a non-expansion over the set of nonnegative functions with domain $\mathcal{S}\times\mathcal{A}$ and $\mathcal{S}$, respectively, when these function spaces are equipped with appropriate norms:
@@
Now, since $f$ is continuous over its domain, by the mean-value theorem, $f$ is decreasing over $[0,\infty)$. ∎
@@
Let $\mu,\nu$ be any distributions over $\mathcal{S}\times\mathcal{A}$ and assume that $\nu$ is admissible. Then, for any $f,f^{\prime}:S\times A\rightarrow[0,\infty)$ we have
@@
Define $(\nu_{i})_{0\leq i\leq k}$ via $\nu_{k}=\nu$ and for $0\leq i<k$, let $\nu_{i}=(\nu_{i+1}P)\times\pi_{f_{i},q^{*}}$. Note that by assumption, $\nu_{k}$ is admissible. It then follows that $\nu_{i}$ for $0\leq i<k$ is also admissible. Indeed, if for some $0\leq i<k$, $\pi=(\pi_{0},\pi_{1},\dots)$ is the nonstationary policy that realizes $\nu_{i+1}$ in step $s\geq 0$, $\pi^{\prime}=(\pi_{0},\pi_{1},\dots,\pi_{s},\pi_{f_{i},q^{*}},\pi_{s+1},\dots)$ is a policy that realizes $\nu_{i}$ in step $s+1$.
@@
Now, we bound $S$ defined above: